OpenStudy (anonymous):

The hypotneuse, c, of right triangle ABC is 5.0 cm long. Given the trignometric ratio cosA=0.75 for angle A, what is the area of the triangle to the nearest tenth of a cm^2

OpenStudy (ddcamp):

\[\cos(a) = 0.75 = \frac{ Adjacent }{ Hypotenuse } = \frac{ Adjacent }{ 5.0 } \rightarrow Adjacent = 3.75\] 3.75² + (Opposite)² = 5.0² Opposite = 3.30718913883 Area = 3.75 * 3.30718913883 = 12.4019592706, which rounds to 12.4

OpenStudy (mertsj):